OpenStudy (anonymous):

evaluate 6x/3y - 4 if x=20 and y=2

OpenStudy (anonymous):

@Directrix

OpenStudy (mysangels):

You would substitute x with 20 and y with 2.

OpenStudy (mysangels):

you would end up with 6(20)/3(2)

OpenStudy (alexandervonhumboldt2):

6*20/6 -4=120/6 -4=20-4=????????

OpenStudy (anonymous):

16

OpenStudy (alexandervonhumboldt2):

correct!

OpenStudy (anonymous):

i have more